OpenStudy (anonymous):

If 5 moles of P4 reacted with 22 moles Cl2 according to the above reaction, determine: a) How many moles PCl3 are produced a)________________ b) How many moles of P4 are left in excess after the reaction (if any) b)________________ c) How many moles of Cl2 are left in excess after the reaction (if any) c)________________

OpenStudy (cwrw238):

P4 + 6Cl2---> 4PCl3 so 1 mole of P4 reacts with 6 Moles of Cl2 to give 4 Moles of PCl3

OpenStudy (cwrw238):

therefore 5 moles of P4 would gives 5*4 = 20 moles of PCl3 and 5*6 = 30 moles of Cl2 would be need to give this result

OpenStudy (cwrw238):

so by proportion number of moles of PCl3 produced is 22 * 20 -- = 14.667 moles 30